In testing the hypotheses

H​0:

p​ = 0.62 vs

Ha​:

p

not equals≠

​0.62, the test statistic is calculated to

be z​ = 2.11. Which of the following is the correct​ p-value?

A.

0.9862

B.

0.0348

C.

0.0174

D.

1.9652
Answer #1

this is two tailed test

z= 2.11

P(z > 2.11 ) = 1 - P(z <  2.11 ) = 1-0.9826=0.0174*2=0.0348

P-value = 0.0348 ( using z table )